A patient with respiratory alkalosis has a blood plasma pH of 7.57. What is the [H3O+] of the blood plasma? Express your answer to two significant figures and include the appropriate units.
Answer #1

As we know, pH = -log[r3 ot] of CH307) = 10 -PH 2) [+20+]=10 -7.57 21 CH2ot) 2.7X1078 M